OpenStudy (anonymous):

2xy+4x+8y+16 how do you factor this out completely??

OpenStudy (tommo_lcfc):

Take 2 out as a common factor to get \[2(xy + 2x + 4y + 8)\] You can also take x as a common factor in the first two terms and 4 out as a common factor in the last two terms to get \[2(x(y + 2) + 4(y + 2))\] This then gives \[2(x + 4)(y + 2)^2\]. Don't quote me on that!

OpenStudy (anonymous):

lol gotcha

OpenStudy (callisto):

starting from second step, it is wrong 2xy+4x+8y+16 = 2(xy +2x +4y+8) = 2[x(y+2) +4(y+2)] = 2(x+4)(y+2)

OpenStudy (anonymous):

kk thanks again lol its much appreciated

OpenStudy (tommo_lcfc):

I told you not to quote me on that answer! :) I thought it wasn't right!

OpenStudy (callisto):

i didn't say your name, you told the others only

OpenStudy (tommo_lcfc):

Sorry :(

OpenStudy (anonymous):

its totally fine lol no worries we all make mistakes you dont even have to and your willingly helpin me out so im not complainin

OpenStudy (anonymous):

start step 1 (2xy+4x) + (8y+16) pair them up.. 2 take common in both like x in first and 4 in the second 3.. x(2y+4) + 4( 2y+4) now take only one bracket of (2y+4) and you will get | (x+4) (2y+4) ....

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you are welcome.